OpenStudy (anonymous):

Solve this 3 step equation plz. .2(d-6) = .3d+5-3+.1d

OpenStudy (anonymous):

Tomhue I love in a non gay way

OpenStudy (anonymous):

Distribute the 0.2 on the left hand side. At the same time, we can combine like terms on the right side. 0.2(d-6)=0.2d-1.2 0.3d+5-3+0.1d=0.4d+2 0.2d-1.2=0.4d+2 Now, we can isolate to find d. 0.2d-1.2=0.4d+2 0.2d-1.2+1.2=0.4d+2+1.2 0.2d=0.4d+3.2 0.2d-0.4d=0.4d-0.4d+3.2 -0.2d=3.2 (those steps could be combined into one on paper. in text form, it would just make everything more confusing.) -0.2d=3.2 -0.2d/-0.2=3.2/-0.2 d=-16

OpenStudy (doc.brown):

Do you mean\[0.2(d-6)=0.3d+5-3+0.1d\]

OpenStudy (anonymous):

I swear to go that's the answer i got as well I just wanted to check it